Question 5 [2;2]


Investigate whether the following functions are odd or even.


(a) 𝑓(π‘₯) = π‘₯^


3


(b) 𝑓(π‘₯) = cos π‘₯

1
Expert's answer
2022-03-10T08:41:06-0500

A function is even if "f(-x)=f(x)"


A function is odd if "f(-x)=-f(x)"


1.

"f(x)=x\u00b3"


"f(-x)=(-x)\u00b3=-x\u00b3=-f(x)"


Hence, "f(x)=x\u00b3" is an odd function


2.

"f(x)=cosx"


"f(-x)=cos(-x)=cosx=f(x)"


Hence, "f(x)=cosx" is an even function